2x + 2y = 38
Y = x + 3
Solve each system using substitution

Answer by ewatrrr(24785)

 
Hi
2x + 2y = 38 and y = (x + 3)
2x +2(x+3) = 38
4x +6 = 48
4x = 42
x = 10.5